# Epistemic Uncertainty in Conformal Scores: A Unified Approach

### v1

@@ 1 Introduction
Machine learning models traditionally focus on point predictions, estimating target variables from input features. However, understanding prediction uncertainty is crucial in many applications [Horta et al., 2015, Freeman et al., 2017, Izbicki and Lee, 2017, Schmidt et al., 2020, Dalmasso et al., 2020, Csillag et al., 2023, Mian et al., 2024, Valle et al., 2024a, Fröhlich et al., 2024]. This has led to increased interest in uncertainty quantification methods, particularly conformal prediction, which constructs predictive regions with finite-sample validity under mild i.i.d. assumptions [Vovk et al., 2005, Shafer and Vovk, 2008]. Unlike probabilistic models that rely on asymptotic assumptions or priors, conformal methods provide a distribution-free framework with guaranteed coverage.
@@
Conformal prediction works by designing a non-conformity score, $s(\mathbf{x},y)$, which measures the degree to which a given label value $y$ aligns with the feature values $\mathbf{x}$ of an instance. Given a new input $\mathbf{x}_{\text{new}}$, the method constructs a predictive region by inverting the non-conformity score at a given confidence level (see Section 2.1). The choice of $s(\mathbf{x},y)$ is critical, as it directly influences the shape and informativeness of the resulting predictive regions [Angelopoulos and Bates, 2021]. For instance, in regression problems, a standard choice is $s(\mathbf{x},y)=|y-g(\mathbf{x})|$, where $g(\mathbf{x})$ is a point prediction for $Y$, typically an estimate of the regression function $\mathbb{E}[Y|\mathbf{x}]$ [Lei et al., 2018]. Another common option is $s(\mathbf{x},y)=\max\{\widehat{q}_{\alpha_{1}}(\mathbf{x})-y,y-\widehat{q}_{\alpha_{2}}(\mathbf{x})\}$, where $\widehat{q}_{\alpha_{1}},\widehat{q}_{\alpha_{2}}$ are quantile estimates of $Y|\mathbf{x}$ [Romano et al., 2019].
@@
Despite offering distribution-free guarantees, standard conformal scores primarily capture aleatoric uncertainty, which arises from inherent randomness in the data generation process - specifically, the fact that $\mathbf{x}$ does not uniquely determine $y$. For example, in the cases discussed above, both $E[Y|\mathbf{x}]$ and $q_{\alpha}(\mathbf{x})$ reflect this form of uncertainty. However, an equally important source of uncertainty is epistemic uncertainty, which stems from limitations in training data and the resulting lack of knowledge about the true data-generating process [Hüllermeier and Waegeman, 2021].
@@
Since standard conformal prediction does not explicitly account for epistemic uncertainty, it can produce misleadingly narrow predictive intervals in regions with little or no training data. Figure 1 illustrates this issue in a regression problem (see Appendix A.1 for technical details): in the range $x\in(7,8)$, there is essentially no training data, so we expect predictive regions to widen, reflecting the increased uncertainty. However, standard conformal methods (e.g., regression-split) instead produce overconfident, narrow intervals in this region. To address this limitation, we propose a novel approach that augments any conformal score with a measure of epistemic uncertainty. As shown in the figure, our method (EPICSCORE) successfully expands the predictive regions where data is scarce, providing a more reliable uncertainty quantification.
@@ 1.1 Novelty
Despite EPICSCORE using Bayesian models, it preserves the finite-sample validity guarantees of conformal prediction and also achieves asymptotic conditional coverage, a property that many existing conformal approaches lack. Furthermore, EPICSCORE is model-agnostic and can be applied on top of any existing conformal score, making it an enhancement rather than a replacement.
@@ 1.2 Relation to Other Work
Several recent frameworks have attempted to model epistemic uncertainty in machine learning predictions (see, e.g., He et al. 2023, Tyralis and Papacharalampous 2024, Wang et al. 2025, Izbicki 2025 and references therein). However, these methods generally lack guarantees on the coverage of their estimates.
@@
Since exact conditional coverage is unattainable without strong assumptions [Lei and Wasserman, 2014], research has focused on two strategies: (1) locally tuning cutoffs to adapt to the data distribution [Boström and Johansson, 2020, Foygel Barber et al., 2021, Guan, 2023, Cabezas et al., 2024, 2025], and (2) designing conformal scores that achieve conditional coverage asymptotically [Romano et al., 2019, Izbicki et al., 2020, Chernozhukov et al., 2021, Izbicki et al., 2022, Dheur et al., 2024, Plassier et al., 2024]. Among these, the score introduced by Dheur et al. [2025, Eq. 14] is particularly relevant to our approach. This method transforms a nonconformity score $s$ via its estimated cumulative distribution function, $s^{\prime}(\mathbf{x};y)=\widehat{F}(s(\mathbf{x},y)|\mathbf{x})$, improving conditional coverage. While $s^{\prime}$ ensures asymptotic conditional coverage, it does not explicitly model epistemic uncertainty. Our approach builds on this idea by incorporating epistemic uncertainty in $s^{\prime}$. In particular, we show that as the calibration sample grows, EPICSCORE achieves asymptotic conditional coverage (Theorem 2).
Bayesian methods have been explored as a means to incorporate epistemic uncertainty into conformal prediction. One approach is to use Bayesian predictive sets and subsequently apply conformal methods to adjust their marginal coverage. However, existing techniques typically do not leverage existing conformal scores explicitly, and do not lead to asymptotic conditional coverage [Vovk et al., 2005, Fong and Holmes, 2021, Wasserman, 2011].
@@
Other efforts to model epistemic uncertainty include Jaber et al. [2024] and Pion and Vazquez [2024], which integrate conformal methods with Gaussian processes. Although these approaches yield calibrated sets for Gaussian processes, they are tailored to this class of models and cannot be applied to other frameworks, such as Bayesian Additive Regression Trees [Chipman et al., 2012]. In contrast, EPICSCORE is model-agnostic and can be applied to any Bayesian model for epistemic uncertainty.
@@ 2.2 Our Approach - EPICSCORE
To construct EPICSCORE, we adopt a Bayesian model that places a prior distribution over $\Theta$ (or equivalently, over $\mathcal{F}$). For simplicity, we assume this prior has a density $f(\theta)$, though the method is generally applicable. This prior captures epistemic uncertainty in the data-generating process. In our experiments, we use Gaussian Processes, Bayesian Additive Regression Trees, and approximate Bayesian Mixture Density Neural Networks with Monte Carlo Dropout, but our approach supports any prior process.
@@
EPICSCORE is summarized in Algorithm 1 and illustrated in Figure 3.
@@ 2.2.1 Special Cases
[Vovk et al., 2005, Valle et al., 2023, 2024b]. Since $Y$ is discrete, the score $s(\mathbf{x},Y)$ is also discrete. Moreover, the cumulative distribution function of its predictive distribution can be computed using the predictive distribution of the labels, $\mathbb{P}(y|\mathbf{x},D)$. In particular, EPICSCORE is given by
@@
  • Both scoring functions in Eqs. 2 and 3 lead to the same EPICSCORE score. Additionally, EPICSCORE follows a similar structure to APS, with the key difference being that the estimates ^$\widehat{\mathbb{P}}(y|\mathbf{x})$ on the sum of Eq. 2 are replaced by the predictive distribution $\mathbb{P}(y|\mathbf{x},D)$.
  • When $\mathbf{x}_{n+1}$ is located in a region with little training data, the estimated probabilities ^$\widehat{\mathbb{P}}(y|\mathbf{x})$ tend to be low for all labels, reflecting a high degree of uncertainty. Consequently, even the most conformal labels will have small probabilities, leading to lower values of $s^{\prime}(\mathbf{x}_{n+1},y)$. This results in larger prediction sets $\{y:s^{\prime}(\mathbf{x}_{n+1},y)\leq t_{1-\alpha}\}$ in areas with sparse data, effectively capturing uncertainty in underrepresented regions.
@@ 3 Theory
Just like any split-conformal method, EPICSCORE guarantees marginal coverage as long as the data is i.i.d., regardless of the chosen Bayesian model (see Appendix D for all proofs):
@@
Assuming that the data are independent and identically distributed (i.i.d.), the confidence region constructed by EPICSCORE  satisfies marginal coverage, that is,
@@
Under Assumption 1, and assuming that the data are independent and identically distributed (i.i.d.), the confidence region constructed by EPICSCORE  satisfies the asymptotic conditional coverage condition, that is:
@@ 4 Experiments
In this section, we evaluate our framework against state-of-the-art baselines, applying EPICSCORE to two initial conformal scores: (i) quantile-based and (ii) regression-based. Each version of EPICSCORE is compared to appropriate baselines, which are detailed in the following subsections.
We consider three versions of EPICSCORE, each using a different model for the predictive distribution (see implementation details in Appendix C):
@@
    where $N(\cdot)$ denotes the normal density and $\pi_{k}(\cdot)$, $\mu_{k}(\cdot)$, $\sigma_{k}(\cdot)$ are all modeled by neural networks, with $\sum_{k=1}^{K}\pi_{k}(\mathbf{x})=1$. To derive a predictive distribution for the scores, we incorporate dropout at each MDN layer. By performing multiple stochastic forward passes using MC Dropout, we approximate the posterior distribution of the MDN parameters, thus propagating uncertainty into the predictive score distribution [Gal and Ghahramani, 2016].
@@
We use the Average Interval Score Loss (AISL) [Gneiting and Raftery, 2007] as our primary evaluation metric, as it balances coverage and interval length, favoring the narrowest valid prediction intervals. Additionally, we assess each method’s efficiency by measuring (i) the average interval length, (ii) the marginal coverage, and (iii) the Pearson correlation between coverage and interval length on Appendix 4, with the latter serving as a proxy for the quality of conditional coverage [Feldman et al., 2021]. A detailed description of all evaluation metrics is provided in Appendix B.1.
@@ 4.1 Quantile-Regression Baselines
For quantile regression-based scores, we adopt CatBoost [Dorogush et al., 2018] as the base quantile-regression model in all conformal methods. See Appendix B.3 for details on hyperparameters.
We compare EPICSCORE to the following baselines:
@@
  • UACQR-P and UACQR-S [Rossellini et al., 2024], which aims to integrate epistemic uncertainty into prediction intervals through ensemble-based statistics. We use their default strategies, deriving UACQR-S correction factors from the ensemble standard deviation and computing UACQR-P percentiles using ensemble order statistics.

### v2

@@ 1 Introduction
Machine learning models traditionally focus on point predictions, estimating target variables from input features. However, understanding prediction uncertainty is crucial in many applications [Horta et al., 2015, Freeman et al., 2017, Izbicki and Lee, 2017, McAllister et al., 2017, Schmidt et al., 2020, Selvan et al., 2020, Vazquez and Facelli, 2022, Csillag et al., 2023, Mian et al., 2024, Valle et al., 2024, Fröhlich et al., 2025]. This has led to increased interest in uncertainty quantification methods, particularly conformal prediction, which constructs predictive regions with finite-sample validity under mild i.i.d. assumptions [Vovk et al., 2005, Shafer and Vovk, 2008]. Unlike probabilistic models that rely on asymptotic assumptions or priors, conformal methods provide a distribution-free framework with guaranteed coverage.
@@
Conformal prediction works by designing a non-conformity score, $s(\mathbf{x},y)$, which measures the degree to which a given label value $y$ aligns with the feature values $\mathbf{x}$ of an instance. Given a new input $\mathbf{x}_{\text{new}}$, the method constructs a predictive region by inverting the non-conformity score at a given confidence level (see Section 2.1). The choice of $s(\mathbf{x},y)$ is critical, as it directly influences the shape and informativeness of the resulting predictive regions [Angelopoulos and Bates, 2021, Izbicki, 2025]. For instance, in regression problems, a standard choice is $s(\mathbf{x},y)=|y-g(\mathbf{x})|$, where $g(\mathbf{x})$ is a point prediction for $Y$, typically an estimate of the regression function $\mathbb{E}[Y|\mathbf{x}]$ [Lei et al., 2018]. Another common option is $s(\mathbf{x},y)=\max\{\widehat{q}_{\alpha_{1}}(\mathbf{x})-y,y-\widehat{q}_{\alpha_{2}}(\mathbf{x})\}$, where $\widehat{q}_{\alpha_{1}},\widehat{q}_{\alpha_{2}}$ are quantile estimates of $Y|\mathbf{x}$ [Romano et al., 2019].
@@
Despite offering distribution-free guarantees, standard conformal scores only capture aleatoric uncertainty, which arises from inherent randomness in the data generation process - specifically, the fact that $\mathbf{x}$ does not uniquely determine $y$. For example, in the cases discussed above, both $E[Y|\mathbf{x}]$ and $q_{\alpha}(\mathbf{x})$ reflect this form of uncertainty. However, an equally important source of uncertainty is epistemic uncertainty, which stems from limitations in training data and the resulting lack of knowledge about the true data-generating process [Hüllermeier and Waegeman, 2021].
@@
Although conformal sets do take epistemic uncertainty into account [Angelopoulos and Bates, 2021] , the conformal scores themselves typically do not. It follows that conformal inference often leads to misleadingly narrow predictive intervals in regions with little or no training data. Figure 1 illustrates this issue in a regression problem (see Appendix A.1 for technical details): in the range $x\in(7,8)$, there is essentially no training data, so we expect predictive regions to widen, reflecting the increased uncertainty. However, standard conformal methods (e.g., regression-split) instead produce overconfident, narrow intervals in this region. To address this limitation, we propose a novel approach that augments any conformal score with a measure of epistemic uncertainty. As shown in the figure, our method (EPICSCORE) successfully expands the predictive regions where data is scarce, providing a more reliable uncertainty quantification.
@@ 1.1 Novelty
Despite EPICSCORE using Bayesian models, it preserves the finite-sample validity guarantees of conformal prediction and also achieves asymptotic conditional coverage, a property that many existing conformal approaches lack. Furthermore, EPICSCORE  can be applied on top of any existing conformal score, making it an enhancement rather than a replacement.
@@ 1.2 Relation to Other Work
Several recent frameworks have attempted to model epistemic uncertainty in machine learning predictions (see, e.g., He et al. 2023, Tyralis and Papacharalampous 2024, Wang et al. 2025 and references therein). However, these methods generally lack guarantees on the coverage of their estimates.
@@
Since exact conditional coverage is unattainable without strong assumptions [Lei and Wasserman, 2014], research has focused on two strategies: (1) locally tuning cutoffs to adapt to the data distribution [Boström and Johansson, 2020, Foygel Barber et al., 2021, Guan, 2023, Cabezas et al., 2024, 2025], and (2) designing conformal scores that achieve conditional coverage asymptotically [Romano et al., 2019, Izbicki et al., 2020, Chernozhukov et al., 2021, Izbicki et al., 2022, Plassier et al., 2024]. Among these, the score introduced by Dheur et al. [2025, Eq. 14] is particularly relevant to our approach. This method transforms a nonconformity score $s$ via its estimated cumulative distribution function, $s^{\prime}(\mathbf{x};y)=\widehat{F}(s(\mathbf{x},y)|\mathbf{x})$, improving conditional coverage. While $s^{\prime}$ ensures asymptotic conditional coverage, it does not explicitly model epistemic uncertainty. Our approach builds on this idea by incorporating epistemic uncertainty in $s^{\prime}$. In particular, we show that as the calibration sample grows, EPICSCORE achieves asymptotic conditional coverage (Theorem 2).
As in EPICSCORE, Bayesian ideas are often used in non-conformal frequentist inference, including what Wasserman [2011] terms “Frasian inference”. Similar ideas appear in I. J. Good’s (Good 1992 and others) advocacy for a Bayes/non-Bayes compromise, in recent work on likelihood-free frequentist inference [Masserano et al., 2023, Dalmasso et al., 2024] and Bayes-optimal sets with frequentist coverage control [Hoff, 2023]. While EPICSCORE’s base model is Bayesian, the final conformal regions remain distribution-free and retain their frequentist distribution–free coverage guarantees, whichever the prior is.
Bayesian methods have also been explored as a means to incorporate epistemic uncertainty into conformal prediction. One approach is to use Bayesian predictive sets and subsequently apply conformal methods to adjust their marginal coverage. However, existing techniques typically do not leverage existing conformal scores explicitly, and do not lead to asymptotic conditional coverage [Vovk et al., 2005, Wasserman, 2011, Fong and Holmes, 2021].
@@
Other efforts to model epistemic uncertainty include Jaber et al. [2024] and Pion and Vazquez [2024], which integrate conformal methods with Gaussian processes. Although these approaches yield calibrated sets for Gaussian processes, they are tailored to this class of models and cannot be applied to other frameworks, such as Bayesian Additive Regression Trees [Chipman et al., 2012]. In contrast, EPICSCORE is model-agnostic (in the sense that it can be used on top of any nonconformity score) and can be applied to any Bayesian model for epistemic uncertainty.
@@ 2.2 Our Approach - EPICSCORE
To construct EPICSCORE, we adopt a Bayesian model that places a prior distribution over $\Theta$ (or equivalently, over $\mathcal{F}$). For simplicity, we assume this prior has a density $f(\theta)$, though the method is generally applicable. The choice of prior plays a crucial role: it can be tailored to capture epistemic uncertainty in the data-generating process (via uninformative or weakly informative priors) or to introduce stronger regularization into the conformal score distribution (via more concentrated priors). In this work, we focus on enhancing the representation of epistemic uncertainty and therefore adopt non-informative priors. However, a balance between capturing uncertainty and enforcing regularization can be explored through prior specification (see Appendix D.5 for a discussion on this trade-off).
In our experiments, we use Gaussian Processes (GP), Bayesian Additive Regression Trees (BART), and approximate Bayesian Mixture Density Networks (MDNs) with Monte Carlo dropout. However, EPICSCORE is compatible with any Bayesian modeling approach that yields a posterior predictive distribution. These three models were chosen for their complementary strengths: BART offers flexibility and interpretability, naturally handling categorical features; GPs provide expressive nonparametric inference with scalable tractability via variational approximations; and MDNs with Monte Carlo dropout are highly scalable, making them well-suited for large-scale or high-dimensional problems.
@@
The posterior $f(\theta|D)$ is not computed explicitly. Instead, Eq. 1 is approximated either via Monte Carlo sampling (as with BART and MC-Dropout) or through a variational approximation of $f(\theta|D)$ (as in the variational GP).
@@
EPICSCORE is summarized in Algorithm 1 and illustrated in Figure 3. Additional details on prior specification and strategies for managing its influence are provided in Appendix D.5. A discussion of the computational complexity introduced by incorporating Bayesian models into our framework is presented in Appendix D.6.
@@ 2.2.1 Special Cases
[Vovk et al., 2005, Valle et al., 2023]. Since $Y$ is discrete, the score $s(\mathbf{x},Y)$ is also discrete. Moreover, the cumulative distribution function of its predictive distribution can be computed using the predictive distribution of the labels, $\mathbb{P}(y|\mathbf{x},D)$. In particular, EPICSCORE is given by
@@
  • Both scoring functions in Eqs. 3 and 4 lead to the same EPICSCORE score. Additionally, EPICSCORE follows a similar structure to APS, with the key difference being that the estimates ^$\widehat{\mathbb{P}}(y|\mathbf{x})$ on the sum of Eq. 3 are replaced by the predictive distribution $\mathbb{P}(y|\mathbf{x},D)$.
@@ 3 Theory
Just like any split-conformal method, EPICSCORE guarantees marginal coverage as long as the data are exchangeable, regardless of the chosen Bayesian model (see Appendix E for all proofs):
@@
Assuming that the data are exchangeable, the confidence region constructed by EPICSCORE  satisfies marginal coverage, that is,
@@
Under Assumption 1, and assuming that the data are exchangeable, the confidence region constructed by EPICSCORE  satisfies the asymptotic conditional coverage condition, that is:
@@
Assumption 1 plays a key role in Theorem 2. While strong, it serves as a practical condition to ensure that the predictive distribution $F(s\mid\mathbf{x},D)$ converges to the true distribution $F(s\mid\mathbf{x},\theta^{*})$ as the sample grows, so that the estimated quantiles approximate the ideal ones. Although potentially stronger than necessary, relaxing it would require a more technical analysis, which we leave for future work. Empirical results in Section 4 show that EPICSCORE achieves coverage close to the nominal level, suggesting robustness even when the assumption holds only approximately.
It is worth emphasizing that the type of guarantee provided by Theorem 2 is, in fact, the strongest achievable in this setting. As shown by Lei and Wasserman [2014], exact conditional coverage is unattainable in general conformal frameworks. Recent methods aim to improve local coverage [Boström and Johansson, 2020, Foygel Barber et al., 2021, Cabezas et al., 2025] and could be combined with EPICSCORE to improve local calibration.
@@ 4 Experiments
In this section, we evaluate our framework against state-of-the-art baselines, applying EPICSCORE to two initial conformal scores: (i) quantile-based and (ii) regression-based. Each version of EPICSCORE is compared to appropriate baselines, which are detailed in the following subsections. In Appendix D.7 we also show that EPICSCORE can be applied to density-based scores.
We consider three versions of EPICSCORE, each using a different model for the predictive distribution (see implementation details in Appendix D):
@@
    where $N(\cdot)$ denotes the normal density and $\pi_{k}(\cdot)$, $\mu_{k}(\cdot)$, $\sigma_{k}(\cdot)$ are all modeled by neural networks, with $\sum_{k=1}^{K}\pi_{k}(\mathbf{x})=1$. To derive a predictive distribution for the scores, we incorporate dropout at each MDN layer. By performing multiple stochastic forward passes using MC Dropout, we approximate the posterior distribution of the MDN parameters, thus propagating uncertainty into the predictive score distribution [Gal and Ghahramani, 2016]. Appendix B explains how MC Dropout yields predictive distributions.
@@
We use the Average Interval Score Loss (AISL) [Gneiting and Raftery, 2007] as our main evaluation metric, as it balances coverage and interval length, promoting narrower yet valid prediction intervals. To evaluate how each method captures epistemic uncertainty, we also report the outlier-to-inlier interval length ratio and average coverage on outliers, which measure adaptivity in data-scarce regions. Additionally, we include (i) average interval length, (ii) marginal coverage, and (iii) the Pearson correlation between coverage and interval length, which serves as a proxy for conditional coverage quality [Feldman et al., 2021]. A full description of all metrics can be found in Appendix C.1, with extended empirical results presented in Appendix 5.
@@ 4.1 Quantile-Regression Baselines
For quantile regression-based scores, we adopt CatBoost [Dorogush et al., 2018] as the base quantile-regression model in all conformal methods. See Appendix C.3 for details on hyperparameters. We compare EPICSCORE to the following baselines:
@@
  • UACQR-P and UACQR-S [Rossellini et al., 2024], that integrate epistemic uncertainty through ensemble-based statistics.We use their default strategies: UACQR-S with ensemble standard deviation, and UACQR-P with ensemble order statistics.
@@
Table 3: Average outlier-to-inlier interval length ratio across methods and datasets in the quantile regression setting. This table reports the average ratio between prediction interval lengths for outliers and inliers across 50 runs. Bold entries denote the best-performing method within a 95% confidence interval. Overall, EPICSCORE consistently achieves higher ratios, demonstrating its effectiveness in adjusting interval widths based on data sparsity.
Dataset EPIC-BART EPIC-GP EPIC-MDN CQR CQR-r UACQR-P UACQR-S airfoil 1.028 (0.034) 1.01 (0.018) 1.048 (0.04) 1.005 (0.013) 1.004 (0.012) 1.002 (0.01) 1.005 (0.016) bike 0.991 (0.007) 1.01 (0.007) 0.989 (0.007) 0.985 (0.007) 0.985 (0.007) 0.99 (0.005) 0.981 (0.01) concrete 0.983 (0.033) 0.995 (0.016) 1.043 (0.029) 0.998 (0.012) 0.998 (0.012) 0.998 (0.008) 0.998 (0.012) cycle 0.953 (0.014) 0.966 (0.014) 0.949 (0.014) 0.997 (0.005) 0.997 (0.005) 1.0 (0.004) 0.998 (0.007) electric 1.01 (0.004) 1.01 (0.004) 1.012 (0.008) 1.008 (0.004) 1.007 (0.003) 1.006 (0.003) 1.01 (0.005) homes 1.116 (0.019) 1.101 (0.016) 1.173 (0.027) 1.088 (0.015) 1.081 (0.014) 1.039 (0.006) 1.094 (0.014) meps19 1.052 (0.029) 1.042 (0.025) 1.044 (0.028) 1.04 (0.025) 1.04 (0.025) 1.021 (0.015) 1.045 (0.026) protein 1.001 (0.002) 1.001 (0.003) 1.006 (0.002) 1.002 (0.001) 1.002 (0.001) 1.001 (0.001) 1.002 (0.001) star 0.994 (0.006) 0.994 (0.003) 0.999 (0.007) 0.994 (0.003) 0.993 (0.003) 0.997 (0.003) 0.997 (0.003) superconductivity 1.021 (0.007) 0.987 (0.006) 1.029 (0.008) 1.041 (0.007) 1.041 (0.007) 1.023 (0.004) 1.049 (0.008) WEC 1.012 (0.006) 1.003 (0.008) 1.02 (0.01) 1.016 (0.003) 1.016 (0.003) 1.011 (0.002) 1.021 (0.004) winered 1.008 (0.012) 1.007 (0.008) 1.072 (0.022) 0.993 (0.008) 0.992 (0.008) 0.996 (0.005) 0.992 (0.009) winewhite 1.031 (0.007) 1.024 (0.006) 1.053 (0.014) 1.018 (0.007) 1.018 (0.007) 1.018 (0.006) 1.034 (0.01)